# Enhancing Privacy in the Early Detection of Sexual Predators Through Federated Learning and Differential Privacy

### v1

@@ 5 Evaluation
Although fine-tuning BERT (Devlin et al. 2019) has been shown to give better results for eSPD (Vogt et al. 2021) (see Appendix C), we use the pre-trained feature-based approach with logistic regression (LR) for our experiments since it is far less computationally expensive and better suited for scaling federated training to a large number of clients. In our framework, each user uses the $\textrm{BERT}_{\textrm{BASE}}$ model to create a context-aware representation of their personal conversation by extracting fixed features from the pre-trained model. The [CLS] representation of the last layer is then used as an input for LR with a binary cross entropy loss function. For each user’s segment, we obtain a $768$ length vector.
6 Experimental Set-Up
FL Implementation.
Warm-up data creation.
Client initialization.
Number of segments
Label OG user non-OG user
0 $10$ $21$
1 $18$ (±$(\pm{14})$ $10$
Training with FL with metric DP
Training with FL with DP-SGD
Training with FL with DP-FedAvg.
eSPD inference.
@@ 7 Empirical Results and Discussion
In this section, we investigate three research questions.
@@ 7.3 RQ3: What are the benefits and shortcomings of different privacy implementations for eSPD?
We first consider a scenario where each user adds noise to their private data. As such, even during local training, the service provider never has access to the raw data. Table 3 shows that the FL model with metric DP has a 52% F-latency with an $\eta=20$. Even with our oversampling technique during training, injecting noise directly into the training data comes with a high utility cost. In Figure 2 in Appendix 6, we can see that models trained with more restrictive privacy budget (e.g. $\eta=10$) are non-performing (e.g. with an F-latency score of 6%).
@@ B.2 The Inference Phase
B.3 The Deployed System
@@
Appendix C Comparison with the literature

### v2

@@ 5 Evaluation
Although fine-tuning BERT (Devlin et al. 2019) has been shown to give better results for eSPD (Vogt et al. 2021) (see Appendix E), we use the pre-trained feature-based approach with logistic regression (LR) for our experiments since it is far less computationally expensive and better suited for scaling federated training to a large number of clients. In our framework, each user uses the $\textrm{BERT}_{\textrm{BASE}}$ model to create a context-aware representation of their personal conversation by extracting fixed features from the pre-trained model. The [CLS] representation of the last layer is then used as an input for LR with a binary cross entropy loss function. For each user’s segment, we obtain a $768$ length vector.
@@ 6 Empirical Results and Discussion
In this section, we investigate three research questions. Details about our implementation can be found in Appendix C, and Appendix D presents additional quantitative results. Our eSPD implementation is available online https://github.com/khaoulachehbouni/fl-espd.
@@ 6.3 RQ3: What are the benefits and shortcomings of different privacy implementations for eSPD?
We first consider a scenario where each user adds noise to their private data. As such, even during local training, the service provider never has access to the raw data. Table 2 shows that the FL model with metric DP has a 52% F-latency with an $\eta=20$. Even with our oversampling technique during training, injecting noise directly into the training data comes with a high utility cost. In Figure 6 in Appendix C, we can see that models trained with more restrictive privacy budget (e.g. $\eta=10$) are non-performing (e.g. with an F-latency score of 6%).
@@ B.2 The Inference Phase
B.3 The Privacy Implementations
Figure 4: Illustration of the different privacy implementations: a lock representing DP is applied first to the embedding representation, then to the local training, and finally to the global training.
B.4 The Deployed System
@@
Appendix C Experimental Set-Up
FL Implementation.
Warm-up data creation.
Client initialization.
Number of segments
Label OG user non-OG user
0 $10$ $21$
1 $18$ (±$(\pm{14})$ $10$
Training with FL with metric DP
Training with FL with DP-SGD
Training with FL with DP-FedAvg.
eSPD inference.
Appendix D Empirical Results
In this section, we explore additional quantitative results.
D.1 Warning Latency
Vogt et al. (2021) define the warning latency as the number of messages exchanged before a warning is raised. In Figure 7, we can see the distribution of the warning latencies for the four private models. We notice that while the FL, DP-SGD and DP-FedAvg models seem to have a similar distribution, for the LDP model, the number of messages it takes to get a final classification is on average higher. At the same time, the maximum of messages reach is also 3 times higher (1531) than the maximum needed for the other models (around 600).
Figure 7: Distribution of the warning latencies for the full predatory conversations for each of our private models.
D.2 Example of Conversation
Similarly, when looking at the difference in speed between models at the conversation level, we notice differences between the different models. For example, the following predatory conversation is flagged from the get-go by the centralized model after the following messages:
  • Predator: “What’s up little girl?”
  • Girl: “Nuttin just chattin”
  • Predator: “Kool”
  • Predator: “So how old are you”
  • Girl: “12/f/nva”
  • Girl: “U?”
While the FL model and the FL model with DPSGD only flag it a dozen messages later, despite exchanges like “are you not going to talk to me cause i’m older than you” and only flags it after “so what you do this past summer”. While the DP-FedAvg model flags it on the first message and the FL model with metric DP only raises a warning much later. This is consistent with the distribution of the warning latencies observed in Figure 7.
D.3 False Positive Evaluation
In this paper, we stated that a 1% FPR strikes a good balance between ensuring system efficiency and avoiding unintended discrimination. In Figure 8, we can see how different false positive rates impact the F1-score and speed of our federated learning model.
Figure 8: Change in Utility and Speed of Detection for different False Positive Rates
Models with higher FPR are faster but also lose in utility. Depending on the area where the model is deployed and the quality of the classifier, different thresholds might be more appropriate.
Appendix E Comparison with the literature